OpenStudy (anonymous):

EASY ALGEBRA 1 QUESTION--> PLEASE BE QUICK FOR EXPLINATIONS, THX!!! :D The table below shows the outputs, y, for different inputs, x: Input (x) : 1 1 3 3 Output (y) : 4 2 12 6 PART A: Does the table represent y as a function of x? Justify your answer. PART B: The total cost f(x), in dollars, for renting a paddle boat for x hours is shown below: f(x) = 20 + 10x What is the value of f(120), and what does f(120) represent?

OpenStudy (caozeyuan):

no, it failed the vertical line test

OpenStudy (anonymous):

Part A) The table doesn't represent y as a function of x. Why? One reason is that 1 or 3 is the first element in more than one ordered pair. Part B) f(x) = 10x + 20 f(120)= 10*120 + 20 = 1200 + 20 = 1220 f(120) represents cost of boating for 120 hours.

OpenStudy (anonymous):

I wasn't very sure of what part a was asking me. I do know that a function can't have a repeating domain, though. Thx for clarifying that one, but how should I justify my answer?

OpenStudy (caozeyuan):

Given a value of x, there cannot be more than one value of y that correspond to x

OpenStudy (anonymous):

Yes